OpenStudy (anonymous):

Match the expressions with the correct LCM 14x3, 15 y z3, & 12y3

OpenStudy (anonymous):

here are the answer choices 420x3y3 z3 70x2 y2 12x2 y2 x2 + 6x + 8 x2 - 9 42xy3 z

OpenStudy (anonymous):

ignoring the unknowns in your expression, do you know how to get the LCM of numbers

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@arch1 @myko

OpenStudy (anonymous):

420x3y3 z3

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

wat about 10x2 , 35x2 y2 ? @myko

OpenStudy (anonymous):

ignoring the unknowns in your expression, do you know how to get the LCM of numbers?

OpenStudy (anonymous):

no but that nots the right answer

OpenStudy (anonymous):

hmm..... that's the right quaestion

OpenStudy (anonymous):

yes and i gave you the answer choices ^^

OpenStudy (anonymous):

is it 70x^2y^2

OpenStudy (anonymous):

i mean that you should learn how to do it, so don't need somebody to just give you answers to same type of questions again and again

OpenStudy (anonymous):

if you whant i explain you...

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

you got two numbers 10 and 35. Put them in prime factor form: 10 = 2*5 35= 7*5 so there are 2,5,7 prime factors involved. To find the LCD multiplie all commun factors. In this case there is just one lonly 5. So LCD =5. To find LCM multiplie all factors comun and not commun. So LCM=2*5*7 = 70 What for x^2y^2 just multiply all uknowns, becouse LCM have to have all of them. So the answer will be : 70x^2y^2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

hmmm i actually understand that a lot!

OpenStudy (anonymous):

ohk let me see if i can do this one alone and i will tell you my answer x -3 , & x2 - 9

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i got 27 but tht is not an answer choice

OpenStudy (anonymous):

just one little addon on my explanation. Might happen that the prime factor form of a nummber is: \[24=2^{3}*3\] so where i said: "To find LCM multiplie all factors comun and not commun" should better say: multiply all factors commun and not commun taken with their highest exponents.

OpenStudy (anonymous):

x -3 , & x2 - 9 is bit different. x2 - 9=(x-3)(x+3) think of (x-3), (x+3) like being prime factors of a simple nummber. The rest is same

OpenStudy (anonymous):

x-3=x-3 x2 - 9=(x-3)(x+3) so factors involved are (x-3) and (x+3) LCM rule says, multiply them, so (x-3)(x+3)=x2 - 9 is LCM

OpenStudy (anonymous):

so its the same thing? x^2-9?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yay!!

OpenStudy (anonymous):

i think you got it, :)

OpenStudy (anonymous):

yes i do!